### v1

# TimeLDM: Latent Diffusion Model for Unconditional Time Series Generation

Abstract
Time series generation is a crucial research topic in the area of deep learning, which can be used for data augmentation, imputing missing values, and forecasting. Currently, latent diffusion models are ascending to the forefront of generative modeling for many important data representations. Being the most pivotal in the computer vision domain, latent diffusion models have also recently attracted interest in other communities, including NLP, Speech, and Geometric Space. In this work, we propose TimeLDM, a novel latent diffusion model for high-quality time series generation. TimeLDM is composed of a variational autoencoder that encodes time series into an informative and smoothed latent content and a latent diffusion model operating in the latent space to generate latent information. We evaluate the ability of our method to generate synthetic time series with simulated and realistic datasets, benchmark the performance against existing state-of-the-art methods. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series. Sores from Context-FID and Discriminative indicate that TimeLDM consistently and significantly outperforms current state-of-the-art benchmarks with an average improvement of 3.4$\times$ and 3.8$\times$, respectively. Further studies demonstrate that our method presents better performance on different lengths of time series data generation. To the best of our knowledge, this is the first study to explore the potential of the latent diffusion model for unconditional time series generation and establish a new baseline for synthetic time series.
[…]
123
1 Introduction
In recent years, time series generation has served a pivotal role in a wide range of applications, such as forecasting, imputing missing values, data augmentation, and simulation in unseen scenarios. Moreover, generating time series data can be a valuable approach to solving the complex challenges associated with data privacy concerns. It allows organizations to unlock the value of their data assets while maintaining the highest standards of privacy protection and regulatory compliance. Numerous studies have used various architectures of deep neural networks for synthetic realistic time series data, including Variational Autoencoder (VAE) based methods [19, 15], Generative Adversarial Network (GAN) based methods [13, 14, 16, 17], Flow-based methods [18] and more [28, 29]. Typically, GAN-based methods have gained plenty of attention from researchers. For instance, Cot-GAN [17] incorporates a specialized loss function based on a regularized Sinkhorn distance, which originates from the principles of causal optimal transport theory. TimeGAN [13] merges the adaptability of the unsupervised GAN framework with the control afforded by supervised training in autoregressive models. However, despite the potential application, the generation results are still unsatisfactory from simulated and realistic time series data.
Figure 1: Distribution of our TimeLDM and Diffusion-TS on multivariate datasets. The blue solid line is for original data, and the yellow dotted line is for synthetic data. TimeLDM aligns better with the original data.
Most recently, diffusion models (DMs) [1, 2] have shown impressive performance on image generation [41, 40] and beyond [39, 38]. DMs construct a diffusion process that gradually degrades the training data structure, and learned neural networks to reverse this contaminated data by progressive denoising. To further improve the synthesis quality of time series data, several studies have also applied DMs for time series synthesis, such as DiffTime [3] adopts future mixup and autoregressive initialization as a condition to generate time information. Diffusion-TS [4] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models. Those methods have emerged as a superior learning architecture to GANs in generative modeling. However, existing approaches often apply diffusion models directly in the data space, which typically consists of limited information and noisy features. Therefore, we are interested in searching for a more flexible framework for time series modeling.
Figure 2: Structure of our proposed TimeLDM. (a) shows the components of TimeLDM, consisting of transformer encoder, reparameterization, diffusion process, reverse process, and transformer decoder. (b) shows the detail of transformer encoder and reparameterization. (c) shows the architecture of latent diffusion model. (d) shows the detail of transformer decoder.
Latent space generation, an efficient alternative model in generative architectures, adopts a pre-trained autoencoder to transfer the generation tasks from the input space to a greater flexible latent domain. In this paper, inspired by the success of the diffusion model on latent space [21, 23], we propose an efficiently synthesize time series method to overcome the above limitations by adopting a smoother and informative latent presentation, named TimeLDM (Time Latent Diffusion Model). As shown in Figure 2 (a), We first transform the raw time series data into an embedding space and train the encoder and decoder network for the VAE. The well-studied VAE converts the time series data into the smoothed and informative latent space. After training the VAE, we apply the encoder to extract the latent vectors from the training set and apply them as the target of the latent diffusion model, which is designed with a denoising MLP. During inference, we generate the latent vectors from the latent diffusion model and then apply the VAE decoder to synthesize the time series.
We validate the performance of our proposed approach for different benchmarks, including simulated and realistic time series datasets. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series. As shown in Table 2, the scores from Context-FID and Discriminative indicate that TimeLDM consistently and significantly outperforms current state-of-the-art benchmarks. For example, TimeLDM achieves an average improvement of 3.8$\times$ in the Discriminative score. Furthermore, we demonstrate that our TimeLDM presents better performance on different lengths of time series data generation, Table 4 presents the results.
The main contributions of this paper are summarized as follows:
  • We propose TimeLDM, a latent diffusion-based method that leverages the high-fidelity image synthesis ability into unconditional time series generation. To the best of our knowledge, this is the first work to explore the potential of LDM for unconditional time series generation and establish a new baseline for synthetic time series.
  • We evaluate the ability of our method to synthetic time series with simulated and realistic datasets, benchmark the performance against existing state-of-the-art methods. Empirically, TimeLDM shows better performance than existing time series generation methods both qualitatively and quantitatively. The ablation study presents the proposed loss function of VAE which plays a crucial role in improving the capability of TimeLDM.
  • Furthermore, we evaluate our method with different lengths of time series data, TimeLDM presents better performance on the proposed benchmark datasets compared with current state-of-the-art methods.
2 Related Works
2.1 Time Series Generation
In a broad range of fields, diverse deep generative models have demonstrated their ability to generate high-quality samples, where generating time series stands as a particularly challenging endeavor within limited information and noisy features. Early methods based on Generative Adversarial Networks [12] (GANs) have been extensively investigated for time series generation. The GAN comprises two neural networks that are trained concurrently: a discriminator trained to differentiate between real samples from the training dataset and those that are generated, and a generator that is trained to produce samples that “fool” the discriminator. For example, TimeGAN [13] applies an embedding function and supervised loss to the original GAN for capturing the temporal dynamics of data throughout time. RTSGAN [14] and PSA-GAN [16] generate long univariate time series samples of high quality with self-attention mechanism. Cot-GAN [17] incorporates a specialized loss function based on a regularized Sinkhorn distance, which originates from the principles of causal optimal transport theory. While Variational Autoencoders [20] (VAEs) and Flows [26, 27] , also drawn attention on researchers. For instance, Fourier Flows [18] is proposed as a method based on normalizing flows followed by a chain of spectral filters leading to an exact likelihood optimization. TimeVAE [19] implements an interpretable temporal structure and achieves reasonable results on time series synthesis.
Recent research [5, 3, 4] has been exploring the use of diffusion models [1] to generate time series, developing on the successes of forward and reverse processing in other areas such as images [8], video [9], text [10], and audio [11]. Among them, TimeGrad [5] is a conditional diffusion model that predicts in an autoregressive manner, with the denoising process guided by the hidden state of a recurrent neural network. CSDI [6] uses self-supervised masking to guide the denoising process like image inpainting. SSSD [7] based on the structured state space model. DiffTime [3] approximated the diffusion function based on CSDI where they remove the side information provided as embedding. They also introduced guided DiffTime to handle new constraints such as trend or fixed-value, without re-training. Diffusion-TS [4] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models.
2.2 Generative Modeling in the Latent Space
Although generative models in the data space have achieved significant success, the latest emerging latent generative models have demonstrated several advantages, including more compact and disentangled representations, robustness to noise, and greater flexibility in controlling generated styles. To enable faster training and inference on limited computational resources. Due to the compaction and smoothness of the latent space, latent generative models are preferred for their computational efficiency. For instance, the recent models [21, 23] have achieved great success in image generation as they exhibit better scaling properties and expressivity than the vanilla diffusion models in the data space. The success of the latent diffusion model in image generation has also inspired their applications in video [24], audio [25], tabular [35], and text [37] domains. In this paper, we explore the application of the latent diffusion models for unconditional time series generation tasks.
3 Method
In this section, we formally introduce TimeLDM, a latent diffusion model (LDM) for unconditional time series generation. Figure 2 presents the overall structure of our TimeLDM, which consists of an encoder-decoder module for variational autoencoder (VAE), reparameterization trick for latent information sampling, and the LDM. In Section 3.1, we formulate the time series generation task first. Then, we introduce the details of VAE in Section 3.2. After that, we unveil the LDM from network architecture to mathematical formulation in Section 3.3. Finally, we summarize the training and sampling procedures in Section 3.4.
3.1 Problem Statement
Let $\mathbf{x}_{1:\tau}=\left(x_{1},\ldots,x_{\tau}\right)\in\mathbb{R}^{\tau\times d}$ be the original time series, where $\tau$ denotes time steps, $d$ is the dimension of observed signals. Given the time series dataset $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$, the aim of TimeLDM is to learn parameterized generative model $p_{\theta}(\mathbf{S})$, which can accurately synthesize diverse and realistic time series data $\hat{\mathbf{x}}\in\hat{{\mathbf{S}}}$ without condition.
3.2 Time Series Autoencoding
To overcome the weakness of data domain generation, we are interested in presenting the time series signals $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$ into an informative and smoothed latent space. The latent representation is $\mathbf{L}=\left\{\mathbf{z}\right\}_{i=1}^{N}$, where $\mathbf{z}_{1:\tau}=\left(z_{1},\ldots,z_{\tau}\right)\in\mathbb{R}^{\tau\times m}$ denotes the latent feature and $m$ is the dimension of representation. The framework of VAE is shown in Figure 2 (b) and (d). As we can see, it designs with encoder and decoder module, VAE $=(\mathcal{E}_{\phi}(\mathbf{x}),\mathcal{D}_{\xi}(\mathbf{z}))$, where the encoder $\mathcal{E}_{\phi}$ learn the latent variable $\mathbf{z}=\mathcal{E}_{\phi}(\mathbf{x})$ and the decoder $\mathcal{D}_{\xi}$ decode latent feature $\mathbf{z}$ back to data domain $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$. Unlike the vanilla VAE apply the ELBO loss function to train, here we adopt $\beta$-VAE [32], where the coefficient $\beta$ adaptively balances the reconstruction loss and KL-divergence loss for effective training.
Algorithm 1 Training Algorithm of TimeLDM
Time series data $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$
Encoder $\mathcal{E}_{\phi}$, Decoder $\mathcal{D}_{\xi}$, Denoising Network $\boldsymbol{\epsilon}_{\theta}$
function Train AutoEncoder
     Initialize $\mathcal{E}_{\phi},\mathcal{D}_{\xi}$
     while $\phi,\xi$ have not converged do
         Sample ${\mathbf{x}}\in{{\mathbf{S}}}$
         Get embedding pattern
         Get positional pattern
         $\mu,\sigma\leftarrow\mathcal{E}_{\phi}(\mathbf{x})$
         $\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
[…]
         $\hat{\mathbf{x}}\leftarrow\mathcal{D}_{\xi}\left(\mathbf{z}\right)$
         $\mathcal{L}=\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\beta\mathcal{L}_{\text{KL}}({\mu},{\sigma})$
         $\phi,\xi\leftarrow\operatorname{optimizer}\left(\mathcal{L};\phi,\xi\right)$
         if $\ell_{\text{recon}}$ fails to decrease for $S$ steps then
              $\beta\leftarrow\lambda\beta$
         end if
     end while
     return $\mathcal{E}_{\phi},\mathcal{D}_{\xi}$
end function
function Train Latent Diffusion
     Initialize $\epsilon_{\theta}$
     while $\theta$ have not converged do
         $\mathbf{z}\sim q_{\phi}\left(\mathbf{z}\mid\mathbf{x}\right)$
         $t\sim\mathbf{U}(0,T)$
         $\varepsilon\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)$
         $\mathbf{z}_{t}=\mathbf{z}_{0}+\varepsilon$
         $\ell(\theta)=\left\|\epsilon_{\theta}\left(\mathbf{z}_{t},t\right)-\varepsilon\right\|_{2}^{2}$
         $\theta\leftarrow\operatorname{optimizer}\left(\mathcal{L}_{LDM};\theta\right)$
     end while
     return $\epsilon_{\theta}$
end function
$\mathcal{E}_{\phi},\mathcal{D}_{\xi}\leftarrow$ Train AutoEncoder
Fix parameters $\phi$ and $\xi$
$\epsilon_{\theta}\leftarrow$ Train Latent Diffusion
return $\mathcal{E}_{\phi},\mathcal{D}_{\xi},\epsilon_{\theta}$
VAE’s Encoder. As shown in Figure 2 (b), the VAE’s Encoder $\mathcal{E}_{\phi}$ first apply one dimension convolutional neural network to learn an embedding pattern $\mathbf{e}=emb(\mathbf{x})\in\mathbb{R}^{\tau\times m}$ from the temporal structures $\mathbf{x}_{1:\tau}\in\mathbb{R}^{\tau\times d}$, then a learnable positional encoding $pe\in\mathbb{R}^{\tau\times m}$ equip to the embedding feature for adaptively learning the time series positional information. After that, we train two transformer encoders to learn the mean $\mu\in\mathbb{R}^{\tau\times m}$ and log variance $\sigma\in\mathbb{R}^{\tau\times m}$ from the positional encoding feature $\mathbf{e}^{pe}_{1:\tau}=\mathbf{e}_{1:\tau}+pe$, respectively. Next, we obtain the latent variables $\mathbf{z}_{1:\tau}$ from the reparameterization trick Function 1.
$\mathbf{z}=\mu+\sigma\cdot\varepsilon,\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (1)
VAE’s Decoder. As shown in Figure 2 (d), the aim of VAE’s Decoder $\mathcal{D}_{\xi}$ is to minimize the reconstruction error by generating outputs that closely resemble the original time series information. The input to the Decoder $\mathcal{D}_{\xi}$ consists of latent variables sampled from a typically Gaussian distribution, which is derived from the Encoder using the reparameterization trick. The architect of VAE decoder is designed with a transformer decoder that incorporates both self-attention and cross-attention mechanisms. The input also respects the learning embedding and positional encoding process. Finally, it generates the realistic samples of time series data $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$.
Training Loss. The training objective of the VAE consists of the reconstruction loss $\mathcal{L}_{\text{recon}}$ and the KL divergence $\mathcal{L}_{\text{KL}}$. The reconstruction loss is composed of the $\mathcal{L}_{1}$ norm and $\mathcal{L}_{2}$ norm in the data domain, as well as the $\mathcal{L}_{1}$ norm in the frequency domain [42] between the input data and the reconstructed signal. Inspire by [29] accurate reconstruction time series signals with Fourier-based loss term, we denote Fast Fourier Transformation loss term $\|\mathcal{FF}\mathcal{T}(\mathbf{x}),\mathcal{FF}\mathcal{T}(\hat{\mathbf{x}})\|$ to learn frequency domain information of time series. $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weights to balance three losses.
$\mathcal{L}_{\text{recon}}=\lambda_{1}\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2}+\lambda_{2}\|\mathbf{x}-\hat{\mathbf{x}}\|+\lambda_{3}\|\mathcal{FFT}\left(\mathbf{x}\right),\mathcal{FFT}\left(\hat{\mathbf{x}}\right)\|$
KL divergence loss regularizes the mean and log variance of the latent space. As shown below 4, the $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ is probabilistic output from the encoder $\mathcal{E}_{\phi}$ that represents the approximate posterior of latent variable $\mathbf{z}$ given the input $\mathbf{x}$; $\mathcal{N}(\mathbf{z};\mu,\sigma)$ is the prior on $\mathbf{z}$ The $\beta$ is adaptively tuned during training, where $\beta=\lambda\beta,\lambda<1$. If the $\mathcal{L}_{\text{recon}}$ fails to decrease with defined steps, the $\beta$ will decrease to encourage the model to pay more attention to the reconstruction term $\mathcal{L}_{\text{recon}}$.
$\mathcal{L}_{\text{K L}}=\beta\mathrm{KL}\left(q_{\boldsymbol{\phi}}(\boldsymbol{\mathbf{z}}\mid\mathbf{x})\|\mathcal{N}(\mathbf{z};\mu,{\sigma})\right)$ (3)
Finally, the overall training objective of the VAE is as below. For the adaptive $\beta$, we set $\beta_{\max}=0.01$, $\beta_{\min}=10^{-5}$, and $\lambda=0.7$, where the $\beta_{\max}$ is initial setting, and $\beta_{\min}$ is the minimum number of the adaptive $\beta$.
$\mathcal{L}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{K L}}$ (4)
3.3 Latent Diffusion Model
After the preparations mentioned above, a trained VAE allows us to access an informative and smoother latent space $\mathbf{L}=\{\mathbf{z}\}_{i=1}^{N}$. Consequently, we are ready to design a diffusion model to generate the latent space information. Figure 2(c) presents the neural network architecture of LDM. First, we reshape the sampling representation into one dimension before passing through a linear layer. Next, we transform the time step $t$ into sinusoidal embeddings $\boldsymbol{t}_{\text{emb}}$, and added to the $\text{Linear}(\mathbf{z})$ After that, we apply four linear layers to learn the denoising pattern. Finally, we reshape the latent representation back to the input shape. Following [2], we adopt below forward process 5 and reverse process 6 to obtain noising data and learn to reverse back:
$\mathbf{z}_{t}=\mathbf{z}_{0}+\sigma(t)\cdot{\varepsilon},\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (5)
$\mathrm{d}\mathbf{z}_{t}=-2\dot{{\sigma}}(t)\sigma(t)\nabla_{\mathbf{z}_{t}}\log p\left(\mathbf{z}_{t}\right)\mathrm{d}t+\sqrt{2\dot{\sigma}(t)\sigma(t)}\mathrm{d}\mathbf{\omega}_{t}$ (6)
where $\mathbf{z}_{0}=\mathbf{z}$ is the original latent representation from encoder, $\mathbf{z}_{t}$ is diffused representation with noise level $\sigma(t)$. While for reverse process, $\nabla_{\mathbf{z}_{t}}\log p_{t}\left(\mathbf{z}_{t}\right)$ preset score of the $\mathbf{z}_{t}$, $\omega_{t}$ is the standard Wiener process. The training object of LDM is:
$\left.\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\mathbf{z}_{0}\sim p\left(\mathbf{z}_{0}\right)}\mathbb{E}_{t\sim p(t)}\mathbb{E}_{{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\|\mathbf{\epsilon}_{\theta}\left(\mathbf{z}_{t},t\right)-\mathbf{\varepsilon}\right\|_{2}^{2}$ (7)
where $\boldsymbol{\epsilon}_{\theta}$ is the neural network to project $\mathbf{z}_{t}$ into Gaussian noise. Following [33], we set the noise level $\sigma(t)=t$, and $\nabla_{\mathbf{z}_{t}}\log p\left(\mathbf{z}_{t}\right)=-\mathbf{\epsilon}_{\theta}\left(\mathbf{z}_{t},t\right)/\sigma(t)$.
Algorithm 2 Sampling Algorithm of TimeLDM
Decoder network $\mathcal{D}_{\xi}$, denoising network $\boldsymbol{\epsilon}_{\theta}$
$\hat{\mathbf{x}}\in\hat{{\mathbf{S}}}$
Sample $\mathbf{z}_{T}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}(T)\mathbf{I}\right),t_{\max}=T$
for $i=\text{max},\dots,1$ do
     $\nabla_{\mathbf{z}_{t_{i}}}\log p(\mathbf{z}_{t_{i}})=-\mathbf{\epsilon}_{\theta}(\mathbf{z}_{t_{i}},t_{i})/\sigma(t_{i})$
     Get $\mathbf{z}_{t_{i-1}}$ via solving the reverse process
end for
$\hat{\mathbf{x}}\sim p_{\xi}\left(\mathbf{x}\mid\mathbf{z}\right)$
return $\hat{{\mathbf{S}}}$
3.4 Training and Sampling
With the proposed formulation and practical parameterization, we now introduce the training and sampling schemes for TimeLDM. The training process of TimeLDM can be divided into two steps where the first step is to train the VAE and the second step is to study the LDM on the latent space. The Algorithm 1 presents the overall training procedure. For sampling process includes generative diffusion data on the standard latent states, reversing the original time series with a well-learning decoder. The Algorithm 2 shows the overall sampling process.
4 Experiments
In this section, we empirically examine the capability of our model for time series generation tasks. We first introduce our experimental setup in Section 4.1. Then, we report and analyze the evaluation main results in Section 4.2. Moreover, we present a further study of TimeLDM on different lengths of time series. Finally, we perform the ablation study of TimeLDM in Section 4.3.
4.1 Experimental Setups
Datasets. As Table 1 shows, we utilize four different datasets to evaluate our model. The datasets include: Sine is a simulated dataset, we generate 10000 samples with 5 features in sinusoidal sequence, and each feature has independent frequencies and phases; Stocks is the Google stock price information from 2004 to 2019, this stock data is presented daily information and includes 6 features. ETTh is built from electricity transformers on 15 minutes basis, including load and oil temperature from July 2016 to July 2018; MuJoCo is the multivariate physics simulation time series data with 14 features.
Table 1: The details of experimental datasets, where the Sines and Mujoco are simulated datasets, the Stocks and ETTh are realistic datasets.
Baseline. We compare our TimeLDM against six unconditional time series generation methods, including: Diffusion-TS [4] (ICLR2024), DiffTime [3] (NeurIPS2023), and Diffwave [11] (ICLR2021), which are based on diffusion architectures. Furthermore, we select TimeGAN [13] (NeurIPS2019), which relies on GAN-based models as same as Cot-GAN [17] (NeurIPS2020). Lastly, we consider TimeVAE [19] (ICLR2022), which is built on a VAE-based approach.
Metric Methods Sines Stocks ETTh MuJoCo
TimeLDM 0.004$\pm$.001 0.032$\pm$.007 0.034$\pm$.003 0.012$\pm$.001
Diffusion-TS 0.006$\pm$.000 0.147$\pm$.025 0.116$\pm$.010 0.013$\pm$.001
Context-FID TimeGAN 0.101$\pm$.014 0.103$\pm$.013 0.300$\pm$.013 0.563$\pm$.052
Score TimeVAE 0.307$\pm$.060 0.215$\pm$.035 0.805$\pm$.186 0.251$\pm$.015
Diffwave 0.014$\pm$.002 0.232$\pm$.032 0.873$\pm$.061 0.393$\pm$.041
DiffTime 0.006$\pm$.001 0.236$\pm$.074 0.299$\pm$.044 0.188$\pm$.028
(Lower the Better) Cot-GAN 1.337$\pm$.068 0.408$\pm$.086 0.980$\pm$.071 1.094$\pm$.079
TimeLDM 0.013$\pm$.005 0.028$\pm$.009 0.028$\pm$.009 0.181$\pm$.029
Diffusion-TS 0.015$\pm$.004 0.004$\pm$.001 0.049$\pm$.008 0.193$\pm$.027
Correlational TimeGAN 0.045$\pm$.010 0.063$\pm$.005 0.210$\pm$.006 0.886$\pm$.039
Score TimeVAE 0.131$\pm$.010 0.095$\pm$.008 0.111$\pm$.020 0.388$\pm$.041
Diffwave 0.022$\pm$.005 0.030$\pm$.020 0.175$\pm$.006 0.579$\pm$.018
DiffTime 0.017$\pm$.004 0.006$\pm$.002 0.067$\pm$.005 0.218$\pm$.031
(Lower the Better) Cot-GAN 0.049$\pm$.010 0.087$\pm$.004 0.249$\pm$.009 1.042$\pm$.007
TimeLDM 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Diffusion-TS 0.006$\pm$.007 0.067$\pm$.015 0.061$\pm$.009 0.008$\pm$.002
Discriminative TimeGAN 0.011$\pm$.008 0.102$\pm$.021 0.114$\pm$.055 0.238$\pm$.068
Score TimeVAE 0.041$\pm$.044 0.145$\pm$.120 0.209$\pm$.058 0.230$\pm$.102
Diffwave 0.017$\pm$.008 0.232$\pm$.061 0.190$\pm$.008 0.203$\pm$.096
DiffTime 0.013$\pm$.006 0.097$\pm$.016 0.100$\pm$.007 0.154$\pm$.045
(Lower the Better) Cot-GAN 0.254$\pm$.137 0.230$\pm$.016 0.325$\pm$.099 0.426$\pm$.022
TimeLDM 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
Diffusion-TS 0.093$\pm$.000 0.036$\pm$.000 0.119$\pm$.002 0.007$\pm$.000
TimeGAN 0.093$\pm$.019 0.038$\pm$.001 0.124$\pm$.001 0.025$\pm$.003
Predictive TimeVAE 0.093$\pm$.000 0.039$\pm$.000 0.126$\pm$.004 0.012$\pm$.002
Score Diffwave 0.093$\pm$.000 0.047$\pm$.000 0.130$\pm$.001 0.013$\pm$.000
DiffTime 0.093$\pm$.000 0.038$\pm$.001 0.121$\pm$.004 0.010$\pm$.001
(Lower the Better) Cot-GAN 0.100$\pm$.000 0.047$\pm$.001 0.129$\pm$.000 0.068$\pm$.009
Original 0.094$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.007$\pm$.001
Table 2: Main results on multiple time series datasets.The best result in each case is bolded.
Setups. In this paper, all the neural networks used for the encoder, latent diffusion, and decoder of TimeLDM are implemented with PyTorch [34] package. For the well-training across all datasets, we tune the limited hyperparameter, as shown in Table 3. Especially, the batch sizes of VAE and LDM are the same. We proceed with the training in two steps. The first step is to train the $\beta$-VAE to obtain latent space information. The second step involves training a diffusion model in the latent space. In the first step, we optimize our network using Adam with default decay rates. The initial learning rate is $10^{-3}$. In the second step, We optimize our network using Adam with $\left(b_{1},b_{2}\right)=(0.9,0.96)$. The initial learning rate start is $10^{-4}$. The main results are training on an NVIDIA RTX 4080 GPU.
Parameter Sines Stocks ETTh MuJoCo
Attention heads 2 2 2 1
Attention head dimension 16 16 16 32
Encoder layers 1 2 2 2
Decoder layers 2 3 3 3
Batch size 1024 512 1024 512
Hidden dimension of LDM 1024 1024 1024 1024
Table 3: Hyperparameters of VAE and latent diffusion model.
Evaluation Methods. For quantitative analysis, we adopt four different evaluation metrics to evaluate the synthesized time series: (1) Context-Fréchet Inception Distance (Context-FID) score [16] assesses the quality of the synthetic time series samples by calculating the difference between representations of time series that fit into the local context; (2) Correlational score [30] assesses temporal dependencies by calculating the absolute error between the cross-correlation matrices of real and synthetic data. (3) Discriminative score [13] evaluates similarity by employing a classification model to differentiate between original and synthetic data in a supervised setting. (4) Predictive score [13] assesses the utility of synthesized data by training a sequence model post-hoc to predict future temporal vectors using the train-synthesis-and-test-real (TSTR) method. For qualitative analysis, we apply two different data representation methods to evaluate the synthesized time series: (1) t-SNE [31] evaluates synthesized time series by projecting both original and synthetic data into a 2-dimensional space; (2) Kernel density estimation is to draw data distributions to check the alignment between original and synthetic data.
4.2 Unconditional Time Series Generation
Main Results. We follow the previous setup in TimeGAN [13] to analyze the performance of models on the benchmark datasets mentioned above. The quantitative evaluation results of 24-length time series generation, which represents the most common basis of comparison in existing works, are listed in Table 2. As can be seen, TimeLDM shows better performance in most cases than Diffusion-TS which is the previous state-of-the-art model. In particular, compared with the Context-FID score and Discriminative score, our TimeLDM achieves state-of-the-art performance on all benchmarks, with the average improvement of 3.4$\times$ in the Context-FID score and 3.8$\times$ in the Discriminative score, respectively. Figure LABEL:fig:ldm shows the qualitative evaluation results of t-SNE, where red dots represent real data instances and blue dots represent generated data samples in all plots. The figure reveals that our methods have better performance with better overlap between the generated data and with original data. Figure 1 presents the distribution alignment of original data and synthetic information. The TimeLDM aligns better with the original data than Diffusion-TS.
Further Analysis. To further confirm the scalability of our TimeLDM, we evaluate the impact of the different time-series lengths on the generative models for unconditional time series. We examine ETTh data with two different lengths, 64 and 128. For these experiments, we keep all the same hyperparameters with the same metrics to assess the generation quality of different methods. The quantitative results are reported in Table 4. As we can see, our proposed TimeLDM can achieve better performance in most evaluation metrics. Especially on the Context-FID score and Discriminative score, TimeLDM realizes significant state-of-the-art performance. The qualitative results are depicted in Figure 3 and 4. Following the previous methods [4], we randomly generate 3000 sequences to project original and synthetic data into a 2-dimensional space. Our TimeLDM shows better alignment than Diffusion-TS with the original data.
Figure 3: Distribution of our TimeLDM and Diffusion-TS on ETTh datasets with different time series lengths. The blue solid line is for original data, and the yellow dotted line is for synthetic data. TimeLDM aligns better with the original ETTh data.
Figure 4: t-SNE visualization on ETTh data. Red is for original data, and blue represents synthetic data. More overlap between the blue and red dots indicates better matching of the data distributions from the generated to the original samples. Our TimeLDM demonstrates a more accurate overlap with the original data across two different lengths of the ETTh dataset.
Metric Methods ETTh-64 ETTh-128
TimeLDM 0.067$\pm$.008 0.169$\pm$.015
Diffusion-TS 0.631$\pm$.058 0.787$\pm$.062
Context-FID TimeGAN 1.130$\pm$.102 1.553$\pm$.169
Score TimeVAE 0.827$\pm$.146 1.062$\pm$.134
Diffwave 1.543$\pm$.153 2.354$\pm$.170
DiffTime 1.279$\pm$.083 2.554$\pm$.318
(Lower the Better) Cot-GAN 3.008$\pm$.277 2.639$\pm$.427
TimeLDM 0.034$\pm$.005 0.058$\pm$.010
Diffusion-TS 0.082$\pm$.005 0.088$\pm$.005
Correlational TimeGAN 0.483$\pm$.019 0.188$\pm$.006
Score TimeVAE 0.067$\pm$.006 0.054$\pm$.007
Diffwave 0.186$\pm$.008 0.203$\pm$.006
DiffTime 0.094$\pm$.010 0.113$\pm$.012
(Lower the Better) Cot-GAN 0.271$\pm$.007 0.176$\pm$.006
TimeLDM 0.030$\pm$.053 0.080$\pm$.044
Diffusion-TS 0.106$\pm$.048 0.144$\pm$.060
Discriminative TimeGAN 0.227$\pm$.078 0.188$\pm$.074
Score TimeVAE 0.171$\pm$.142 0.154$\pm$.087
Diffwave 0.254$\pm$.074 0.274$\pm$.047
DiffTime 0.150$\pm$.003 0.176$\pm$.015
(Lower the Better) Cot-GAN 0.296$\pm$.348 0.451$\pm$.080
TimeLDM 0.115$\pm$.010 0.117$\pm$.009
Diffusion-TS 0.116$\pm$.000 0.110$\pm$.003
TimeGAN 0.132$\pm$.008 0.153$\pm$.014
Predictive TimeVAE 0.118$\pm$.004 0.113$\pm$.005
Score Diffwave 0.133$\pm$.008 0.129$\pm$.003
DiffTime 0.118$\pm$.004 0.120$\pm$.008
(Lower the Better) Cot-GAN 0.135$\pm$.003 0.126$\pm$.001
Table 4: Further evaluaction on long-term ETTh dataset generation. The best result in each case is bolded.
4.3 Ablation Study
In this part, we first assess the adaptive $\beta$ with the fixed values in the VAE model. Then, we analyze the effectiveness of the reconstruction loss function of VAE. We compare the loss function with its three variants: (1) w/o FFT term during training, (2) w/o $\mathcal{L}_{1}$ norm term during training, (3) w/o $\mathcal{L}_{2}$ norm term during training. The ablation study across all the benchmarks presents the results in Table 5 and Table 6, respectively.
The effect of adaptive $\beta$. We evaluate the adaptive weighting coefficient $\beta$ in the VAE model. Table 5 presents the results of adaptive $\beta$ and constant $\beta$ on the aforementioned datasets. As can be seen, the difference in performance between the Sines and Stocks benchmarks is not significant. While there is a significant performance disparity between the ETTh and MuJoCo benchmarks. The adaptive $\beta$ improve the effectiveness of TimeLDM, remarkably. This emphasizes the superior performance demonstrated by the adaptive $\beta$ approach in training the VAE model.
Metric $\beta$ Sines Stocks ETTh MuJoCo
Discriminative Adaptive 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Score $10^{-2}$ 0.004$\pm$.004 0.015$\pm$.015 0.034$\pm$.019 0.095$\pm$.007
(Lower the Better) $10^{-5}$ 0.004$\pm$.004 0.020$\pm$.019 0.093$\pm$.004 0.010$\pm$.007
Adaptive 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
Predictive $10^{-2}$ 0.093$\pm$.000 0.037$\pm$.000 0.122$\pm$.005 0.010$\pm$.001
Score $10^{-5}$ 0.093$\pm$.000 0.038$\pm$.000 0.123$\pm$.005 0.009$\pm$.002
(Lower the Better) Original 0.094$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.007$\pm$.001
Table 5: Ablation study for the adaptive $\beta$, which balances the reconstruction loss and KL loss. The best result in each case is bolded.
The effect of reconstruction loss. We evaluate the effectiveness of reconstruction loss term in the VAE model. Table 6 presents the results of three variants on the aforementioned datasets. As can be seen, the performance gap among them for the Predictive score is negligible, while for the Discriminative score, as the Figure LABEL:fig:ABS, the FFT-loss term shows a crucial role in improving the capability of TimeLDM.
Metric Methods Sines Stocks ETTh MuJoCo
TimeLDM 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Discriminative w/o FFT 0.008$\pm$.003 0.022$\pm$.028 0.013$\pm$.009 0.010$\pm$.009
Score w/o $\mathcal{L}_{1}$ 0.007$\pm$.005 0.020$\pm$.015 0.014$\pm$.014 0.004$\pm$.006
(Lower the Better) w/o $\mathcal{L}_{2}$ 0.006$\pm$.005 0.014$\pm$.010 0.010$\pm$.008 0.005$\pm$.005
TimeLDM 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
w/o FFT 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.006 0.008$\pm$.002
Predictive w/o $\mathcal{L}_{1}$ 0.093$\pm$.000 0.037$\pm$.000 0.121$\pm$.004 0.008$\pm$.001
Score w/o $\mathcal{L}_{2}$ 0.093$\pm$.000 0.037$\pm$.000 0.121$\pm$.006 0.008$\pm$.002
(Lower the Better) Original 0.094$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.007$\pm$.001
Table 6: Ablation study for VAE reconstruction loss function. The best result in each case is bolded.
Metric Methods Sines Stocks ETTh MuJoCo
Discriminative Score w/o VAE 0.019 $\pm$.015 0.062$\pm$0.029 0.011$\pm$.007 0.023$\pm$.028
(Lower the Better) TimeLDM 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Predictive Score w/o VAE 0.095 $\pm$.000 0.039 $\pm$.000 0.124 $\pm$.003 0.011$\pm$.001
(Lower the Better) TimeLDM 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
Table 7: Ablation study for VAE representation. The best result in each case is bolded.
The effect of Latent Space Information. As shown in the Table 7, the w/o VAE method means the removal of Variational Autoencoder(VAE), thereby directly generating time series information in the data space. Obviously, the absence of VAE for mapping feature data into latent space results in significant performance drops.
5 Conclusion
In this paper, we propose TimeLDM, a novel latent diffusion model for unconditional time series generation. Particularly, we explore diffusion on the informative and smoother latent space where the original time series is encoded by a variational autoencoder. We evaluate our method on the simulated and realistic datasets, benchmark the performance against existing state-of-the-art methods. Experimental results demonstrate that TimeLDM persistently delivers high-quality generated data both qualitatively and quantitatively. Scores from Context-FID and Discriminative indicate that TimeLDM surpasses the current state-of-the-art in time-series data generation with an average improvement of 3.4$\times$ and 3.8$\times$, respectively. Further studies demonstrate that our method yields better performance on different lengths of time series data generation. To the best of our knowledge, this is the first work to explore the potential of the latent diffusion model for unconditional time series generation.
{ack}
We would like to thank all the anonymous reviewers for their constructive feedback.