### v2

# TimeLDM: Latent Diffusion Model for Unconditional Time Series Generation

Abstract
Time series generation is a crucial research topic in the area of decision-making systems, which can be particularly important in domains like autonomous driving, healthcare, and, notably, robotics. Recent approaches focus on learning in the data space to model time series information. However, the data space often contains limited observations and noisy features. In this paper, we propose TimeLDM, a novel latent diffusion model for high-quality time series generation. TimeLDM is composed of a variational autoencoder that encodes time series into an informative and smoothed latent content and a latent diffusion model operating in the latent space to generate latent information. We evaluate the ability of our method to generate synthetic time series with simulated and real-world datasets and benchmark the performance against existing state-of-the-art methods. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series. For example, TimeLDM achieves new state-of-the-art results on the simulated benchmarks and an average improvement of 55% in Discriminative score with all benchmarks. Further studies demonstrate that our method yields more robust outcomes across various lengths of time series data generation. Especially, for the Context-FID score and Discriminative score, TimeLDM realizes significant improvements of 80% and 50%, respectively. The code will be released after publication.
I INTRODUCTION
Time series generation holds a pivotal role across numerous applications, such as robotics [1, 2], autonomous driving [3, 4], and healthcare [5, 6]. Additionally, generating time series can be a valuable approach to solving the complex challenges associated with data privacy concerns. It enables agents to learn a wealth of information without containing any actual sensitive data, providing a safer framework for model training and development.
Numerous studies have used various architectures of deep neural networks for synthetic realistic time series data, including Variational Autoencoder (VAE) based methods [7, 8], Generative Adversarial Network (GAN) based methods [9, 10, 11, 12], and Diffusion-based methods [13, 14]. Typically, Diffusion-based methods have gained plenty of attention from researchers. For instance, DiffTime [13] adopts future mix-up and autoregressive initialization as a condition to generate time information. Diffusion-TS [14] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models. Those methods have emerged as a superior learning architecture in generative modeling to others. However, existing approaches often apply learning models directly in the data space, which typically consists of limited information and noisy features. Therefore, we are interested in searching for a more flexible framework for time series modeling.
Figure 1: t-SNE visualization on the stocks dataset, TimeLDM shows better overlap between the generated data and original data than TimeVAE.
Figure 2: Structure of our proposed TimeLDM. (a) shows the components of TimeLDM, consisting of the transformer encoder, reparameterization, diffusion process, reverse process, and transformer decoder. (b) shows the details of the transformer encoder and reparameterization. (c) shows the architecture of the latent diffusion model. (d) shows the details of the transformer decoder.
Latent space generation, an efficient alternative model in generative architectures, adopts a pre-trained autoencoder to transfer the generation tasks from the input space to a greater flexible latent domain. In this paper, inspired by the success of the diffusion model on latent space [15, 16], we propose an efficiently synthesized time series method to overcome the above limitations by adopting a smoother and informative latent presentation, named TimeLDM (Time Latent Diffusion Model). As shown in Figure 2 (a), We first transform the raw time series data into an embedding space and train the encoder and decoder network for the VAE. The well-studied VAE converts the time series data into the latent space. After that, we apply the latent information as the target of the latent diffusion model (LDM), which is designed with a denoising MLP. During inference, we generate the latent vectors from the LDM and then apply the VAE decoder to synthesize the time series.
We validate the performance of our proposed approach for different benchmarks, including simulated and real-world time series datasets. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series (see Figure~5). In Table II and III, the Discriminative scores of TimeLDM consistently outperform current state-of-the-art benchmarks. Furthermore, Table IV demonstrates that TimeLDM presents better performance on different lengths of time series data generation. The main contributions of this paper are summarized as follows:
  • We propose TimeLDM, a latent diffusion-based method that leverages the high-fidelity image synthesis ability into unconditional time series generation. To the best of our knowledge, this is the first work to explore the potential of LDM for unconditional time series generation.
  • We evaluate the ability of our method to synthetic time series with simulated and real-world datasets. Empirically, TimeLDM shows better performance than existing generation methods both qualitatively and quantitatively. The ablation study presents the proposed loss function of VAE, which plays a crucial role in improving the capability of our method.
  • Furthermore, we evaluate TimeLDM with different lengths of time series data, which presents better performance on the proposed benchmark datasets compared with current state-of-the-art methods.
II RELATED WORKS
Time Series Generation. Deep generative models have demonstrated their ability to generate high-quality samples across a wide array of fields, where generating time series stands as a particularly challenging endeavor within limited information and noisy features. Early methods based on GANs [17] have been extensively investigated for time series generation. For example, TimeGAN [9] applies an embedding function and supervised loss to the original GAN for capturing the temporal dynamics of data throughout time. Cot-GAN [12] incorporates a specialized loss function based on a regularized Sinkhorn distance, which originates from the principles of causal optimal transport theory. While VAEs [18] also drew the attention of researchers. For instance, TimeVAE [7] implements an interpretable temporal structure and achieves reasonable results on time series synthesis. Recent research [13, 14] has been exploring the use of diffusion models [19] to generate time series, developing on the successes of forward and reverse processing in other areas such as images [20], video [21], text [22], and audio [23]. Among them, DiffTime [13] approximated the diffusion function based on CSDI [24] where they remove the side information provided as embedding. Diffusion-TS [14] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models.
Generative Modeling in the Latent Space. Although generative models in the data space have achieved significant success, the latest emerging LDMs [15, 16] have demonstrated several advantages, including more compact and disentangled representations, robustness to noise, and greater flexibility in controlling generated styles. LDMs have achieved great success in image generation as they exhibit better scaling properties and expressivity than the vanilla diffusion models in the data space. The success of the LDM in image generation has also inspired their applications in video [25], audio [26], tabular [27], and text [28] domains. In this paper, we explore the application of the LDM for unconditional time series generation tasks.
Algorithm 1 Training Algorithm of TimeLDM
Time series data $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$
Encoder $\mathcal{E}_{\phi}$, Decoder $\mathcal{D}_{\xi}$, Denoising Network $\boldsymbol{\epsilon}_{\theta}$
function Train AutoEncoder
    Initialize $\mathcal{E}_{\phi},\mathcal{D}_{\xi}$
    while $\phi,\xi$ have not converged do
         Sample ${\mathbf{x}}\in{{\mathbf{S}}}$
         Get the embedding pattern
         Get the positional pattern
         $\mu,\sigma\leftarrow\mathcal{E}_{\phi}(\mathbf{x})$
         $\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
[…]
         $\hat{\mathbf{x}}\leftarrow\mathcal{D}_{\xi}\left(\mathbf{z}\right)$
         $\mathcal{L}=\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\beta\mathcal{L}_{\text{KL}}({\mu},{\sigma})$
         $\phi,\xi\leftarrow\operatorname{optimizer}\left(\mathcal{L};\phi,\xi\right)$
         if $\ell_{\text{recon}}$ fails to decrease for $S$ steps then
             $\beta\leftarrow\lambda\beta$
         end if
    end while
    return $\mathcal{E}_{\phi},\mathcal{D}_{\xi}$
end function
function Train Latent Diffusion
    Initialize $\epsilon_{\theta}$
    while $\theta$ have not converged do
         $\mathbf{z}\sim q_{\phi}\left(\mathbf{z}\mid\mathbf{x}\right)$
         $t\sim\mathbf{U}(0,T)$
         $\varepsilon\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)$
         $\mathbf{z}_{t}=\mathbf{z}_{0}+\varepsilon$
         $\ell(\theta)=\left\|\epsilon_{\theta}\left(\mathbf{z}_{t},t\right)-\varepsilon\right\|_{2}^{2}$
         $\theta\leftarrow\operatorname{optimizer}\left(\mathcal{L}_{LDM};\theta\right)$
    end while
    return $\epsilon_{\theta}$
end function
$\mathcal{E}_{\phi},\mathcal{D}_{\xi}\leftarrow$ Train AutoEncoder
Fix parameters $\phi$ and $\xi$
$\epsilon_{\theta}\leftarrow$ Train Latent Diffusion
return $\mathcal{E}_{\phi},\mathcal{D}_{\xi},\epsilon_{\theta}$
III METHOD
TimeLDM, as shown in Figure 2, consists of an encoder-decoder module for VAE, a reparameterization trick for latent information sampling, and the LDM. In this section, we formulate the time series generation task first. Then, we introduce the details of VAE and LDM from network architecture to mathematical formulation. Finally, we summarize the training and sampling procedures.
III-A Problem Statement
Let $\mathbf{x}_{1:\tau}=\left(x_{1},\ldots,x_{\tau}\right)\in\mathbb{R}^{\tau\times d}$ be the original time series, where $\tau$ denotes time steps, $d$ is the dimension of observed signals. Given the time series dataset $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$, the aim of TimeLDM is to learn parameterized generative model $p_{\theta}(\mathbf{S})$, which can accurately synthesize diverse and realistic time series data $\hat{\mathbf{x}}\in\hat{{\mathbf{S}}}$ without condition.
III-B Time Series Autoencoding
To overcome the weakness of data domain generation, we are focusing on presenting the time series signals $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$ into an informative and smoothed latent space. The latent representation is $\mathbf{L}=\left\{\mathbf{z}\right\}_{i=1}^{N}$, where $\mathbf{z}_{1:\tau}=\left(z_{1},\ldots,z_{\tau}\right)\in\mathbb{R}^{\tau\times m}$ denotes the latent feature and $m$ is the dimension of representation. The framework of VAE is shown in Figure 2 (b) and (d). As we can see, it designs with encoder and decoder module, VAE $=(\mathcal{E}_{\phi}(\mathbf{x}),\mathcal{D}_{\xi}(\mathbf{z}))$, where the encoder $\mathcal{E}_{\phi}$ learn the latent variable $\mathbf{z}=\mathcal{E}_{\phi}(\mathbf{x})$ and the decoder $\mathcal{D}_{\xi}$ decode latent feature $\mathbf{z}$ back to data domain $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$. Here we adopt $\beta$-VAE [29], the coefficient $\beta$ adaptively balances the reconstruction loss and KL-divergence loss for effective training.
VAE’s Encoder. As shown in Figure 2 (b), the VAE’s Encoder $\mathcal{E}_{\phi}$ first apply a convolutional neural network to learn an embedding pattern $\mathbf{e}=emb(\mathbf{x})\in\mathbb{R}^{\tau\times m}$ from the temporal structures $\mathbf{x}_{1:\tau}\in\mathbb{R}^{\tau\times d}$, then a learnable positional encoding $pe\in\mathbb{R}^{\tau\times m}$ equip to the embedding feature for adaptively learning the time series positional information. After that, we train two transformer encoders to learn the mean $\mu\in\mathbb{R}^{\tau\times m}$ and log variance $\sigma\in\mathbb{R}^{\tau\times m}$ from the positional encoding feature $\mathbf{e}^{pe}_{1:\tau}=\mathbf{e}_{1:\tau}+pe$, respectively. Next, we obtain the latent variables $\mathbf{z}_{1:\tau}$ from the reparameterization trick Function 1.
$\mathbf{z}=\mu+\sigma\cdot\varepsilon,\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (1)
VAE’s Decoder. As shown in Figure 2 (d), the aim of VAE’s Decoder $\mathcal{D}_{\xi}$ is to minimize the reconstruction error by generating outputs that closely resemble the original time series information. The input to the Decoder $\mathcal{D}_{\xi}$ consists of latent variables sampled from a typically Gaussian distribution, which is derived from the Encoder using the reparameterization trick. The architect of the VAE decoder incorporates both self-attention and cross-attention mechanisms. The input also respects the learning embedding and positional encoding process. Finally, it generates the realistic samples of time series data $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$.
Training Loss. The training objective of the VAE consists of the reconstruction loss $\mathcal{L}_{\text{recon}}$ and the KL divergence $\mathcal{L}_{\text{KL}}$. The reconstruction loss is composed of the $\mathcal{L}_{1}$ norm, $\mathcal{L}_{2}$ norm in the data domain, and the Fast Fourier Transformation (FFT) [30] loss term $\|\mathcal{FF}\mathcal{T}(\mathbf{x}),\mathcal{FF}\mathcal{T}(\hat{\mathbf{x}})\|$ in the frequency domain [31], which is inspired by HyperTime [32] for accurate time series reconstruction. $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weights to balance three losses.
$\mathcal{L}_{\text{recon}}=\lambda_{1}\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2}+\lambda_{2}\|\mathbf{x}-\hat{\mathbf{x}}\|+\lambda_{3}\|\mathcal{FFT}\left(\mathbf{x}\right),\mathcal{FFT}\left(\hat{\mathbf{x}}\right)\|$
KL divergence loss regularizes the mean and log variance of the latent space. As shown Equation 3, the $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ is probabilistic output from the encoder $\mathcal{E}_{\phi}$ that represents the approximate posterior of latent variable $\mathbf{z}$ given the input $\mathbf{x}$; $\mathcal{N}(\mathbf{z};\mu,\sigma)$ is the prior on $\mathbf{z}$ The $\beta$ is adaptively tuned during training, where $\beta=\lambda\beta,\lambda<1$. If the $\mathcal{L}_{\text{recon}}$ fails to decrease with defined steps, the $\beta$ will decrease to encourage the model to pay more attention to the reconstruction term.
$\mathcal{L}_{\text{KL}}=\beta\mathrm{KL}\left(q_{\boldsymbol{\phi}}(\boldsymbol{\mathbf{z}}\mid\mathbf{x})\|\mathcal{N}(\mathbf{z};\mu,{\sigma})\right)$ (3)
Finally, the overall training objective of the VAE is as below. For the adaptive $\beta$, we set $\beta_{\max}=10^{-2}$, $\beta_{\min}=10^{-5}$, and $\lambda=0.7$, where the $\beta_{\max}$ is initial setting, and $\beta_{\min}$ is the minimum number of the adaptive $\beta$.
$\mathcal{L}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{KL}}$ (4)
Algorithm 2 Sampling Algorithm of TimeLDM
Decoder network $\mathcal{D}_{\xi}$, denoising network $\boldsymbol{\epsilon}_{\theta}$
$\hat{\mathbf{x}}\in\hat{{\mathbf{S}}}$
Sample $\mathbf{z}_{T}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}(T)\mathbf{I}\right),t_{\max}=T$
for $i=\text{max},\dots,1$ do
    $\nabla_{\mathbf{z}_{t_{i}}}\log p(\mathbf{z}_{t_{i}})=-\mathbf{\epsilon}_{\theta}(\mathbf{z}_{t_{i}},t_{i})/\sigma(t_{i})$
    Get $\mathbf{z}_{t_{i-1}}$ via solving the reverse process
end for
$\hat{\mathbf{x}}\sim p_{\xi}\left(\mathbf{x}\mid\mathbf{z}\right)$
return $\hat{{\mathbf{S}}}$
III-C Latent Diffusion Model
After the preparations mentioned above, a trained VAE allows us to access the latent space $\mathbf{L}=\{\mathbf{z}\}_{i=1}^{N}$. Figure 2(c) presents the neural network architecture of LDM. First, we reshape the sampling representation into one dimension before passing through a linear layer. Next, we transform the time step $t$ into sinusoidal embeddings $\boldsymbol{t}_{\text{emb}}$, and added to the $\text{Linear}(\mathbf{z})$ After that, we apply four linear layers to learn the denoising pattern. Finally, we reshape the latent representation back to the input shape. Following [33], we adopt below forward process Equation 5 and reverse process Equation 6 to obtain noising data and learn to reverse back:
$\mathbf{z}_{t}=\mathbf{z}_{0}+\sigma(t)\cdot{\varepsilon},\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (5)
$\mathrm{d}\mathbf{z}_{t}=-2\dot{{\sigma}}(t)\sigma(t)\nabla_{\mathbf{z}_{t}}\log p\left(\mathbf{z}_{t}\right)\mathrm{d}t+\sqrt{2\dot{\sigma}(t)\sigma(t)}\mathrm{d}\mathbf{\omega}_{t}$ (6)
where $\mathbf{z}_{0}=\mathbf{z}$ is the original latent representation from encoder, $\mathbf{z}_{t}$ is diffused representation with noise level $\sigma(t)$. While for reverse process, $\nabla_{\mathbf{z}_{t}}\log p_{t}\left(\mathbf{z}_{t}\right)$ preset score of the $\mathbf{z}_{t}$, $\omega_{t}$ is the standard Wiener process. The training object of LDM is:
$\left.\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\mathbf{z}_{0}\sim p\left(\mathbf{z}_{0}\right)}\mathbb{E}_{t\sim p(t)}\mathbb{E}_{{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\|\mathbf{\epsilon}_{\theta}\left(\mathbf{z}_{t},t\right)-\mathbf{\varepsilon}\right\|_{2}^{2}$ (7)
where $\boldsymbol{\epsilon}_{\theta}$ is the neural network to project $\mathbf{z}_{t}$ into Gaussian noise. Following [34], we set the noise level $\sigma(t)=t$, and $\nabla_{\mathbf{z}_{t}}\log p\left(\mathbf{z}_{t}\right)=-\mathbf{\epsilon}_{\theta}\left(\mathbf{z}_{t},t\right)/\sigma(t)$.
Parameter Sines MuJoCo Stocks ETTh fMRI
dim(x) 5 14 6 7 50
Attention heads 2 2 2 2 2
Attention head dimension 16 16 16 16 16
Encoder layers 1 1 2 2 1
Decoder layers 2 2 3 3 2
Batch size 1024 1024 512 1024 1024
Hidden dimension of LDM 1024 4096 1024 1024 4096
TABLE I: Hyperparameters of VAE and LDM.
III-D Training and Sampling
With the proposed formulation and practical parameterization, we now introduce the training and sampling schemes for TimeLDM. The training process of TimeLDM can be divided into two steps where the first step is to train the VAE and the second step is to study the LDM on the latent space. The Algorithm 1 presents the overall training procedure. For sampling process includes generative diffusion data on the standard latent states, reversing the original time series with a well-learning decoder. The Algorithm 2 shows the overall sampling process.
IV EXPERIMENTS
We evaluate TimeLDM for time series generation with five different benchmarks, covering simulated and real-world datasets. Our framework demonstrates better performance than existing methods, both qualitatively and quantitatively. Further analysis across various lengths of time series data confirms the robustness of TimeLDM.
Metric Methods Sines MuJoCo
TimeLDM 0.004$\pm$.001 0.006$\pm$.000
Diffusion-TS 0.006$\pm$.000 0.013$\pm$.001
Context-FID TimeGAN 0.101$\pm$.014 0.563$\pm$.052
Score TimeVAE 0.307$\pm$.060 0.251$\pm$.015
DiffTime 0.006$\pm$.001 0.188$\pm$.028
(Lower the Better) Cot-GAN 1.337$\pm$.068 1.094$\pm$.079
TimeLDM 0.013$\pm$.005 0.189$\pm$.029
Diffusion-TS 0.015$\pm$.004 0.193$\pm$.027
Correlational TimeGAN 0.045$\pm$.010 0.886$\pm$.039
Score TimeVAE 0.131$\pm$.010 0.388$\pm$.041
DiffTime 0.017$\pm$.004 0.218$\pm$.031
(Lower the Better) Cot-GAN 0.049$\pm$.010 1.042$\pm$.007
TimeLDM 0.006$\pm$.005 0.004$\pm$.004
Diffusion-TS 0.006$\pm$.007 0.008$\pm$.002
Discriminative TimeGAN 0.011$\pm$.008 0.238$\pm$.068
Score TimeVAE 0.041$\pm$.044 0.230$\pm$.102
DiffTime 0.013$\pm$.006 0.154$\pm$.045
(Lower the Better) Cot-GAN 0.254$\pm$.137 0.426$\pm$.022
TimeLDM 0.093$\pm$.000 0.007$\pm$.000
Diffusion-TS 0.093$\pm$.000 0.007$\pm$.000
TimeGAN 0.093$\pm$.019 0.025$\pm$.003
Predictive TimeVAE 0.093$\pm$.000 0.012$\pm$.002
Score DiffTime 0.093$\pm$.000 0.010$\pm$.001
(Lower the Better) Cot-GAN 0.100$\pm$.000 0.068$\pm$.009
Original 0.094$\pm$.001 0.007$\pm$.001
TABLE II: Main results on simulated time series datasets. The best result in each case is bolded.
Figure 3: Visualizations of the simulated MuJoCo dataset, synthesized by TimeLDM, Diffusion-TS and TimeVAE.
IV-A Experimental Setups
Datasets. We utilize five different datasets to evaluate our model, including: Sine is a simulated dataset with 5 features in sinusoidal sequence, and each feature has independent frequencies and phases [9]; MuJoCo is the multivariate physics simulation time series data with 14 features [35]; Stocks is the Google stock price information from 2004 to 2019, presented daily information and includes 6 features [7]; ETTh is built from electricity transformers on 15 minutes basis, including load and oil temperature from July 2016 to July 2018 [36]; fMRI serves as a benchmark for causal discovery, featuring simulations that realistically mimic blood-oxygen-level-dependent time series [37].
Baseline. We compare our TimeLDM against five unconditional time series generation methods, including Diffusion-based architectures (Diffusion-TS [14] and DiffTime [13]), GAN-based models (TimeGAN [9] and Cot-GAN [12]), and VAE-based approach (TimeVAE [7]).
Metric Methods Stocks ETTh fMRI
TimeLDM 0.032$\pm$.007 0.034$\pm$.003 0.139$\pm$.025
Diffusion-TS 0.147$\pm$.025 0.116$\pm$.010 0.105$\pm$.006
Context-FID TimeGAN 0.103$\pm$.013 0.300$\pm$.013 1.292$\pm$.218
Score TimeVAE 0.215$\pm$.035 0.805$\pm$.186 14.449$\pm$.969
DiffTime 0.236$\pm$.074 0.299$\pm$.044 0.340$\pm$.015
(Lower the Better) Cot-GAN 0.408$\pm$.086 0.980$\pm$.071 7.813$\pm$.550
TimeLDM 0.028$\pm$.009 0.028$\pm$.009 1.036$\pm$.025
Diffusion-TS 0.004$\pm$.001 0.049$\pm$.008 1.411$\pm$.042
Correlational TimeGAN 0.063$\pm$.005 0.210$\pm$.006 23.502$\pm$.039
Score TimeVAE 0.095$\pm$.008 0.111$\pm$.020 17.296$\pm$.526
DiffTime 0.006$\pm$.002 0.067$\pm$.005 1.501$\pm$.048
(Lower the Better) Cot-GAN 0.087$\pm$.004 0.249$\pm$.009 26.824$\pm$.449
TimeLDM 0.017$\pm$.011 0.009$\pm$.003 0.102$\pm$.020
Diffusion-TS 0.067$\pm$.015 0.061$\pm$.009 0.167$\pm$.023
Discriminative TimeGAN 0.102$\pm$.021 0.114$\pm$.055 0.484$\pm$.042
Score TimeVAE 0.145$\pm$.120 0.209$\pm$.058 0.476$\pm$.044
DiffTime 0.097$\pm$.016 0.100$\pm$.007 0.245$\pm$.051
(Lower the Better) Cot-GAN 0.230$\pm$.016 0.325$\pm$.099 0.492$\pm$.018
TimeLDM 0.037$\pm$.000 0.118$\pm$.007 0.099$\pm$.000
Diffusion-TS 0.036$\pm$.000 0.119$\pm$.002 0.099$\pm$.000
TimeGAN 0.038$\pm$.001 0.124$\pm$.001 0.126$\pm$.002
Predictive TimeVAE 0.039$\pm$.000 0.126$\pm$.004 0.113$\pm$.003
Score DiffTime 0.038$\pm$.001 0.121$\pm$.004 0.100$\pm$.000
(Lower the Better) Cot-GAN 0.047$\pm$.001 0.129$\pm$.000 0.185$\pm$.003
Original 0.036$\pm$.001 0.121$\pm$.005 0.090$\pm$.001
TABLE III: Main results on real-world time series datasets. The best result in each case is bolded.
Setups. In this paper, all the neural networks of TimeLDM are implemented with PyTorch [38] package. For the well-training across all datasets, we tune the limited hyperparameter, as shown in Table I. We proceed with the training in two steps. The first step is to train the $\beta$-VAE to obtain latent space information. The second step involves training a diffusion model in the latent space. In the first step, we optimize our network using Adam with default decay rates. The initial learning rate is $10^{-3}$. In the second step, We optimize our network using the Adam optimizer with the first and second moment decay rates set to $0.9$ and $0.96$ respectively. The initial learning rate start is $10^{-4}$. The main results are training on an NVIDIA RTX 4080 GPU.
Figure 4: Visualizations of the real-world ETTh dataset, synthesized by TimeLDM, Diffusion-TS, and TimeVAE.
Evaluation Methods. For quantitative analysis, we adopt four different evaluation metrics to evaluate the synthesized time series: (1) Context-Fréchet Inception Distance (Context-FID) score assesses the quality of the synthetic time series samples by calculating the difference between representations of time series that fit into the local context [11]; (2) Correlational score assesses temporal dependencies by calculating the absolute error between the cross-correlation matrices of real and synthetic data [39]; (3) Discriminative score evaluates similarity by employing a classification model to differentiate between original and synthetic data in a supervised setting [9]; (4) Predictive score assesses the utility of synthesized data by training a sequence model post-hoc to predict future temporal vectors using the train-synthesis-and-test-real (TSTR) method [9]. For qualitative analysis, we apply two different data representation methods to evaluate the synthesized time series: (1) t-SNE evaluates synthesized time series by projecting both original and synthetic data into a two-dimensional space [40]; (2) Kernel density estimation is to draw data distributions to check the alignment between original and synthetic data.
Figure 5: Examples of generating time series from the fMRI dataset. Our approach yields the closest results to the original training data.
IV-B Unconditional Time Series Generation
Main Results. We follow the previous setup in TimeGAN [9] to analyze the performance of models on the benchmark datasets mentioned above. The quantitative evaluation results of 24-length time series generation, which represents the most common comparison in existing works, are listed in Table II and III. As can be seen, TimeLDM achieves state-of-the-art results on the simulated benchmarks. Compared with the discriminative score, TimeLDM achieves an average improvement of 55% over Diffusion-TS [14] in all benchmarks. Figure 3 and 4 show the qualitative evaluation results of t-SNE and Kernel density. For the t-SNE analysis, where a greater overlap of blue and red dots shows a better distributional similarity between the generated data and original data. Figure 3 and 4 of t-SNE reveals that our methods have better overlap between the generated data and original data. The Kernel density presents the distribution alignment of original data and synthetic information. Based on the figure, our TimeLDM aligns better with the original data than Diffusion-TS and TimeVAE. We also present the generating time series from the fMRI dataset in Figure 5. Compared to the Diffusion-TS [14] and TimeVAE [7], TimeLDM generates time series that more closely resemble the original training set, while TimeVAE [7] struggles to learn features from the fMRI dataset.
Metric Methods ETTh-64 ETTh-128
TimeLDM 0.067$\pm$.008 0.169$\pm$.015
Diffusion-TS 0.631$\pm$.058 0.787$\pm$.062
Context-FID TimeGAN 1.130$\pm$.102 1.553$\pm$.169
Score TimeVAE 0.827$\pm$.146 1.062$\pm$.134
DiffTime 1.279$\pm$.083 2.554$\pm$.318
(Lower the Better) Cot-GAN 3.008$\pm$.277 2.639$\pm$.427
TimeLDM 0.034$\pm$.005 0.058$\pm$.010
Diffusion-TS 0.082$\pm$.005 0.088$\pm$.005
Correlational TimeGAN 0.483$\pm$.019 0.188$\pm$.006
Score TimeVAE 0.067$\pm$.006 0.054$\pm$.007
DiffTime 0.094$\pm$.010 0.113$\pm$.012
(Lower the Better) Cot-GAN 0.271$\pm$.007 0.176$\pm$.006
TimeLDM 0.030$\pm$.053 0.080$\pm$.044
Diffusion-TS 0.106$\pm$.048 0.144$\pm$.060
Discriminative TimeGAN 0.227$\pm$.078 0.188$\pm$.074
Score TimeVAE 0.171$\pm$.142 0.154$\pm$.087
DiffTime 0.150$\pm$.003 0.176$\pm$.015
(Lower the Better) Cot-GAN 0.296$\pm$.348 0.451$\pm$.080
TimeLDM 0.115$\pm$.010 0.117$\pm$.009
Diffusion-TS 0.116$\pm$.000 0.110$\pm$.003
TimeGAN 0.132$\pm$.008 0.153$\pm$.014
Predictive TimeVAE 0.118$\pm$.004 0.113$\pm$.005
Score DiffTime 0.118$\pm$.004 0.120$\pm$.008
(Lower the Better) Cot-GAN 0.135$\pm$.003 0.126$\pm$.001
TABLE IV: Further evaluation on long-term ETTh dataset generation. The best result in each case is bolded.
Figure 6: Visualizations of the real-world ETTh dataset, synthesized by TimeLDM and Diffusion-TS with 64 and 128-time series lengths.
Further Analysis. To further confirm the scalability of our TimeLDM, we evaluate the impact of the different time-series lengths on the generative models for unconditional time series. We examine ETTh data with two different lengths, 64 and 128. For these experiments, we keep all the same hyperparameters with the same metrics to assess the generation quality of different methods. The quantitative results are reported in Table IV. As we can see, our proposed TimeLDM can achieve better performance in most evaluation metrics. Especially on the Context-FID score and Discriminative score, TimeLDM realizes significant state-of-the-art performance with 80% and 50% improvement over Diffusion-TS [14]. The qualitative results are depicted in Figure 5. Our TimeLDM shows better alignment than Diffusion-TS with the original data.
IV-C Ablation Study
In this part, we first assess the adaptive $\beta$ with the fixed values ($\beta_{\max}$, $\beta_{\min}$) in the VAE model. Then, we analyze the effectiveness of the reconstruction loss function of VAE. We compare the loss function with its three variants: (1) w/o FFT loss term during training, (2) w/o $\mathcal{L}_{1}$ norm term during training, (3) w/o $\mathcal{L}_{2}$ norm term during training. The ablation study across all the benchmarks presents the results in Table V and Table VI, respectively.
The effect of adaptive $\beta$. We evaluate the adaptive weighting coefficient $\beta$ in the VAE model. Table V presents the results of adaptive $\beta$ and constant values ($\beta_{\max}$, $\beta_{\min}$) on the aforementioned datasets. As can be seen, the difference in performance between the Sines and Stocks benchmarks is insignificant. At the same time, there is a significant performance disparity between the ETTh and MuJoCo benchmarks. The adaptive $\beta$ improve the effectiveness of TimeLDM, remarkably. This emphasizes the superior performance demonstrated by the adaptive $\beta$ approach in training the VAE model.
Metric $\beta$ Sines MuJoCo Stocks ETTh fMRI
Discriminative Adaptive 0.006$\pm$.005 0.004$\pm$.004 0.017$\pm$.011 0.009$\pm$.003 0.102$\pm$.020
Score $10^{-2}$ 0.004$\pm$.004 0.258$\pm$.023 0.015$\pm$.015 0.034$\pm$.019 0.496$\pm$.003
(Lower the Better) $10^{-5}$ 0.004$\pm$.004 0.007$\pm$.008 0.020$\pm$.019 0.093$\pm$.004 0.216$\pm$.021
Adaptive 0.093$\pm$.000 0.007$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.099$\pm$.000
Predictive $10^{-2}$ 0.093$\pm$.000 0.013$\pm$.001 0.037$\pm$.000 0.122$\pm$.005 0.100$\pm$.000
Score $10^{-5}$ 0.093$\pm$.000 0.007$\pm$.001 0.038$\pm$.000 0.123$\pm$.005 0.100$\pm$.000
(Lower the Better) Original 0.094$\pm$.001 0.007$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.090$\pm$.001
TABLE V: Ablation study for the adaptive $\beta$, which balances the reconstruction loss and KL loss. The best result in each case is bolded.
Metric Methods Sines MuJoCo Stocks ETTh fMRI
TimeLDM 0.006$\pm$.005 0.004$\pm$.004 0.017$\pm$.011 0.009$\pm$.003 0.102$\pm$.020
Discriminative w/o FFT 0.008$\pm$.003 0.005$\pm$.002 0.022$\pm$.028 0.013$\pm$.009 0.115$\pm$.019
Score w/o $\mathcal{L}_{1}$ 0.007$\pm$.005 0.008$\pm$.004 0.020$\pm$.015 0.014$\pm$.014 0.107$\pm$.019
(Lower the Better) w/o $\mathcal{L}_{2}$ 0.006$\pm$.005 0.007$\pm$.006 0.014$\pm$.010 0.010$\pm$.008 0.129$\pm$.014
TimeLDM 0.093$\pm$.000 0.007$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.099$\pm$.000
w/o FFT 0.093$\pm$.000 0.008$\pm$.002 0.037$\pm$.000 0.118$\pm$.006 0.099$\pm$.000
Predictive w/o $\mathcal{L}_{1}$ 0.093$\pm$.000 0.008$\pm$.002 0.037$\pm$.000 0.121$\pm$.004 0.099$\pm$.000
Score w/o $\mathcal{L}_{2}$ 0.093$\pm$.000 0.007$\pm$.001 0.037$\pm$.000 0.121$\pm$.006 0.099$\pm$.000
(Lower the Better) Original 0.094$\pm$.001 0.007$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.090$\pm$.001
TABLE VI: Ablation study for VAE reconstruction loss function. The best result in each case is bolded.
The effect of reconstruction loss. We evaluate the effectiveness of reconstruction loss term in the VAE model. Table VI presents the results of three variants on the aforementioned datasets. As can be seen, the performance gap among them for the Predictive score is negligible, the FFT loss term and $\mathcal{L}_{1}$ norm term show a crucial role in improving the capability of TimeLDM for the Discriminative score.
V CONCLUSIONS
In this paper, we propose TimeLDM, a novel latent diffusion model for unconditional time series generation. Particularly, we explore diffusion on the latent space where the original time series is encoded by a variational autoencoder. We evaluate our method on the simulated and real-world datasets and benchmark the performance against existing state-of-the-art methods. Experimental results demonstrate that TimeLDM persistently delivers high-quality generated data both qualitatively and quantitatively. Remarkably, TimeLDM achieves new state-of-the-art results on the simulated benchmarks and an average improvement of 55% in Discriminative score with all benchmarks. Further studies demonstrate that our method yields better performance on different lengths of time series data generation. To the best of our knowledge, this is the first work to explore the potential of the latent diffusion model for unconditional time series generation. We hope that TimeLDM can serve as a robust baseline for generating informative time series tokens for agents learning in the field of physical AI.